OpenStudy (anonymous):

List the first six terms of the following sequences. a) t1=1, tn= 1/1+tn-1 (n is greater to and equal to 2) b) t1= 1, t2=2, tn=tn-1 - tn-2 (n is greater to and equal to 3)

ganeshie8 (ganeshie8):

\[\large t_1 = 1\] \[\large t_n = \dfrac{1}{1+t_{n-1}}\] plugin \(n=2\)

OpenStudy (anonymous):

1/2?

ganeshie8 (ganeshie8):

Yes! \[\large t_1 = 1\] \[\large t_2 = \dfrac{1}{2}\] \[\large t_n = \dfrac{1}{1+t_{n-1}}\] plugin \(n=3\)

OpenStudy (anonymous):

2/3?

ganeshie8 (ganeshie8):

Yep! keep going..

OpenStudy (anonymous):

oh i see! i think i got it. 3/5, 5/8, 8/13 How do you do b)?

ganeshie8 (ganeshie8):

b isn't any different : \[t_1 = 1\] \[t_2 = 2\] \[t_n = t_{n-1}-t_{n-2}\] plugin \(n=3\)

OpenStudy (anonymous):

1, then I got -1 for n=4

ganeshie8 (ganeshie8):

looks good

OpenStudy (anonymous):

but the answer key says 2

OpenStudy (anonymous):

the pattern is 1,2,1, -1,-2,-1

ganeshie8 (ganeshie8):

that loooks good !

OpenStudy (anonymous):

where did we get the 2?

ganeshie8 (ganeshie8):

1, `2`,1, -1,-2,-1 you're talking about that `2` ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

we didn't get that `2`, it was given to us we're given first TWO terms in the question itself..

ganeshie8 (ganeshie8):

Notice this sequence requires you to take the difference of previous TWO terms so you need TWO starting values as seed for the recursive sequence..

OpenStudy (anonymous):

but for n=5, i got 0, not -2

ganeshie8 (ganeshie8):

look at first four terms : \[ 1,~ 2, ~1, ~ -1, \]

ganeshie8 (ganeshie8):

you can get the next term by taking difference of -1 and 1, yes ? whats the value of \(-1-1\) ?

OpenStudy (anonymous):

-2, opps forgot the neg. sign. thank you!